# Model for Peanuts: Hijacking ML models without Training Access is Possible

### v1

@@ 4. SnatchML: General Approach
(3) $k=\underset{i\in[1,m]}{Argmin}\Big{[}\delta(\zeta_{h}(x^{*}_{i}),\zeta_{h}(x_{s}))\Big{]}$
@@
In essence, SnatchML  hijacks the model at inference time by exploiting its capability to distinguish patterns in a data distribution related to a task $\mathcal{T}^{\prime}$ that has not been included in its training.
@@
Remark. It is worth noting that in the proposed approach, there are no a priori constraints on $m$. In fact, in state-of-the-art attacks, such as (Salem et al., 2022), the number of classes inferred through the hijacking task cannot exceed those in the original task, i.e., $m\leq n$. Therefore, our approach not only assumes a stronger threat model but also overcomes this fundamental limitation, enabling the attacker to hijack an $n$-class model for an $m$-class task where $m>n$. We illustrate this case in Section 5. Since there is no interference with the model during training, the same model can be hijacked to execute more than one task at the same time. This case is illustrated in Section 6.
In the following sections, we illustrate our approach to different scenarios where the attacker tries to misuse a benign trained model to hijack the victim model for a different task. Sections 5, 7, and 6 correspond to the case where the hijacking task and the original task share relevant features. In Section 8, we investigate the general case where the hijacking tasks are unrelated to the original one.
5. Scenario 1: Emotion Recognition
5.1. Context
@@
Let the original task $\mathcal{T}$ be Emotion Recognition (ER). We aim to explore the extent to which a model trained to recognize emotions can be vulnerable to hijacking at inference time for the purpose of biometric identification, denoted as the hijacking task $\mathcal{T^{\prime}}$. In practice, the adversary possesses a dataset of ’persons of interest’ (the hijacking dataset) and intends to identify these individuals by exploiting the original model.
5.2. Setup
Datasets. To run our experiments, we need datasets that are labeled for both the original and the hijacking tasks. For this experiment, we use $4$ different datasets from both real and synthetic distributions:
@@ 5.2. Setup
(i) CK+: The Extended Cohn-Kanade (CK+) (Lucey et al., 2010) is a dataset for emotion recognition that contains $593$ video clips from $123$ individuals between the ages of $18$ and $50$. These videos capture transitions between neutral and peak emotions: anger, disgust, fear, happiness, sadness, and surprise.
@@
(ii) Olivetti faces: The Olivetti faces Dataset (Samaria and Harter, 1994) is a facial recognition dataset comprising 400 grayscale face images gathered from 40 unique individuals with ten images for each. The dataset captures variations in lighting and facial expression with a uniform black background. The dataset is labeled with encoded integers referring to the identities of the 40 individuals.
@@
(iii) Celebrity dataset. We extracted 79 facial images for $9$ different celebrities. Then, we generated emotion-specific images from the neutral samples using BRIA AI (BRIA AI, [n. d.]) is based on a visual generative tool (Elasri et al., 2022), we generated emotion-specific images from the neutral samples. These images are labeled for seven emotions (anger, disgust, fear, happiness, sadness, surprise, and neutral) and associated with $9$ unique individuals used as identity labels.
@@
(iv) Synthetic dataset: To cover the synthetic data distribution case, we create a new dataset labeled for emotions and identities using MakeHuman (Bastioni et al., 2008), an open-source tool designed to generate virtual human faces. This dataset encompasses $395$ images labeled with both emotion and identity. Among these images, $47$ depicts neutral expressions, while $348$ represents six distinct emotions. The dataset contains $47$ individual identities, each associated with varying emotional expressions.
Implementation. We evaluate this attack scenario by designing an experimental setup with four pre-trained ER models: (i) 2D-CNN (Allaert et al., 2022; Poux et al., 2021) with three convolutional layers, each followed by ReLU activation and max-pooling followed by two fully connected layers, (ii) ResNet-9 (He et al., 2016), (iii) MobileNet (Howard et al., 2017), and (iv) Vision Transformer (Dosovitskiy et al., 2020). We train the models for ER on CK+ (Lucey et al., 2010) as the original task $\mathcal{T}$ using a learning rate of $0.001$ and Adam optimizer for $100$ epochs.
Comparison. To evaluate the accuracy of the hijacking tasks, we consider a lower bound and an upper bound (UB): The lower bound (LB) corresponds to the random guessing probability, while the UB corresponds to an unconstrained version of (Salem et al., 2022) without covertness requirement, which is practically equivalent to freely poisoning the model by training it on both the original and hijacking tasks.
@@
Figure 3. Examples of the hijacked ER model’s output top-5 candidates for users re-identification from CK+ dataset.
@@
Figure 4. SnatchMLperformance on hijacking an ER model for users re-identification. ’Rnd’ and ’Hijacking UB’ refer to the random guess and upper bound ratios, respectively.
5.3. ER Systems for user re-identification
We refer to this scenario as user re-identification, as it involves the adversary aiming to identify users whose data (images) were used to train the original ER model. It is worth noticing that in our experiments, we run the hijacking on users’ data that is not member of the original task’s training dataset as it does not comply with our threat model.
@@ 5.3. ER Systems for user re-identification
We use the top-N ranked reference users as an evaluation metric for the hijacking attack performance. The top-N metric is particularly relevant for evaluating identification tasks (Wang and Deng, 2021); it can provide a contextual method for narrowing down potential candidate users. Figure 3 illustrates with queries and their corresponding top-5 samples. In the second line of this figure, the query image features a person with dark skin, and only one individual with dark skin appears in the top-5 reference users. The attacker could reasonably deduce that the unknown query image likely belongs to this individual. Figure 4 shows the top-1 to top-5 accuracy results for the hijacking task, presented as mean and standard deviation after conducting $10$-fold experiments. Considering our hijacking task involves identifying $85$ persons, we use random classification as a reference ($1.1\%$).SnatchMLachieves over $40\%$ top-1 accuracy for 2D-CNN, while the hijacking UB is up to $98\%$.
Although all four tested models demonstrate significant performance, the results depicted in Figure 4 reveal disparities among them, suggesting that different architectures may have varying levels of susceptibility to the proposed model hijacking for the same task. For example, while the attack on 2D-CNN and ResNet-9 succeeded with around $40\%$ accuracy, the ViT and MobileNet were more robust with an average Top-5 accuracy not exceeding $30\%$.
5.4. ER systems learn biometric identification
@@
In this second scenario, we consider a setting where the attacker aims to hijack the model for biometric identification of users whose data is not necessarily a part of the ER model’s training dataset. Following the same approach in Section 5.3, we exploit the pre-trained ER model to identify unseen users based on the extracted BEK. As for the hijacking reference database, we consider two main settings:
@@
Figure 5. ER model used to identify users from the Olivetti dataset (Samaria and Harter, 1994). The top-5 similar users from the hijacking reference database are displayed for each query.
Figure 6. SnatchMLperformance on hijacking an ER model for biometric identification of users from the Olivetti dataset. ’Rnd’ and ’Hijacking UB’ refer to the random guess and upper bound ratios, respectively.
Figure 7. SnatchMLperformance on hijacking an ER model for biometric identification of users from the real (Celebrity) and synthetic datasets. Rnd refers to random guess.
@@
Case (i): We follow the same approach as detailed in Section 5.3 by training the ER model on CK+ dataset. However, for evaluating the identification attack, we use the Olivetti dataset (Samaria and Harter, 1994). We perform ER queries with images from the Olivetti dataset to get the output FV/Logits and execute SnatchML. Figure 5 displays illustration samples of the identification task for Olivetti. An interesting observation is highlighted in the first row of the figure, where the top-5 output candidates for a query involving a person wearing glasses also predominantly feature candidates wearing glasses. This indicates that the ER model has learned to recognize this accessory despite its irrelevance from an emotion recognition perspective. The second row in Figure 5 also highlights the significance of the top-5 metric. In this case, the query image corresponds to a female individual with $4$ out of $5$ of the closest candidates representing male subjects, with the first female candidate appearing only at rank-5, potentially leading to a precise identification.
Figure 6 shows the hijacking attack accuracy results on the four pre-trained ER models regarding top-1 to top-5 accuracy. The attacks demonstrate notable success, with the top-1 accuracy achieving over $60\%$ on 2D-CNN and ResNet-9. Although the lower success rate on MobileNet and the ViT remains consistent in this setting, the hijacking performance is still considerable, considering a random guessing probability of $\sim$2.5%.
@@
Case (ii): In this scenario, we assume that the attacker has access only to neutral images of targeted users. To conduct our attack, we use two datasets featuring individuals with neutral facial expressions, namely Celebrity and Synthetic, as detailed in Section 5.1. The ER model is pretrained on the CK+ dataset. We assess the effectiveness of the identification hijacking attack by performing ER queries with images from both Celebrity and Synthetic datasets, analyzing the output FV/Logits. The accuracy of the hijacking task using the four pretrained ER models is depicted in Figure 7. Notably, the top-1 identification accuracy reached approximately $100\%$ in the 2D-CNN model under both black-box and white-box attacks for both datasets. The least successful attack was observed with MobileNet under the black-box scenario, where it achieved approximately $48\%$ top-1 identification accuracy with the Synthetic dataset while the random guess probability is around $2\%$.
6. Scenario 2: Age/Gender/Ethinicity
@@
Table 1. Accuracy results of SnatchML  on the UTK dataset. We train four (04) benign ML models (2D-CNN, ResNet-9, MobileNet, and ViT) to predict Age, Gender, and Ethnicity. We hijack each model to infer other properties (e.g., from age, we infer gender/ethnicity).
@@
Hijacking Age Accuracy Random 0.166 0.166 0.166 0.166
@@
Hijacking Gender Accuracy Random 0.500 0.500 0.500 0.500
@@
Hijacking Ethnicity Accuracy Random 0.200 0.200 0.200 0.200
@@ 6.1. Context
In this section, we consider an application that predicts personal attributes (e.g., age, gender, ethnicity) from facial images. An example of this application is Microsoft Azure’s Face (Microsoft, [n. d.]) as an MLaaS API, where users can query one specific model (e.g., for age estimation) using facial images. While the original task $\mathcal{T}$ here is age estimation, we assume that an adversary can hijack the same model to infer other personal attributes like gender or ethnicity. In practice, we assume the adversary has access to a publicly available facial image database labeled with personal attributes (relevant to the hijacking task). The implications of this hijacking attack can be critical, mainly if users don’t consent to using their data. Unauthorized inference of gender and ethnicity from this data could constitute a significant privacy violation, potentially leading to discriminatory practices, especially against individuals from marginalized ethnic or gender groups.
6.2. Setup
@@
Dataset: UTKface The UTKFace (Zhang et al., 2017) dataset is a collection of $20,000$ facial images of individuals ranging in age from 1 to 116 years old. Each image is labeled with information on age, gender, and ethnicity as follows:
-Age: (6 Classes) – ’Newborns/Toddlers’, ’Pre-adolescence’, ’Teenagers’, ’Young Adults’, ’Middle-aged’, ’Seniors’.
-Gender: (2 classes) – ’Male’ and ’Female’.
-Ethnicity: (5 Classes) – ’White’, ’Black’, ’Asian’, ’Indian’, ’Others’.
Models. We consider the same model architectures as in Section 5.2, trained using a learning rate of $10^{-3}$ and Adam optimizer. We compare our results with the hijacking upper bound (UB), which corresponds to an unconstrained version of (Salem et al., 2022) without covertness requirement, which is practically equivalent to freely poisoning the model by training it on both the original and hijacking tasks.
6.3. Cross-attribute hijacking: age, gender and ethnicity
@@
The adversary’s goal is to hijack a model – exclusively trained on one of the three tasks– to infer other personal attributes for which the victim model has not been trained. Specifically, from an unknown facial image query submitted by an unknown user to an age estimation model and by only using the original model’s BEK, the adversary aims to infer the gender and ethnicity of the user. For a comprehensive study, we test all the possible combinations of (age, gender, and ethnicity) as original and hijacking tasks. Following the leave-one-out on the possible combinations, we exclusively train a model on one task and test SnatchML  on the two remaining tasks. Following our threat model, we run SnatchML  using the facial images from the test dataset.
@@
Table 1 provides an overview of the experimental results with random guess probability as a lower bound. We expectedly note that the white-box setting consistently delivers higher performance. For instance, ResNet-9, trained for gender recognition, can be hijacked to predict age and ethnicity with an accuracy of approximately $45\%$, while the random guessing probability for this task is $16.60\%$ and the upper bound hijacking accuracy is $66.89\%$. Different tasks have varying levels of difficulty that can be related to their inherent complexity. For example, age recognition is challenging; a 17-year-old ’teenager’ and an 18-year-old ’young adult’ are not easily distinguished. The fact that models trained for age prediction as the original task achieve accuracy ranging from $63\%$ (Transformer) to $70\%$ (ResNet-9) illustrates the task complexity and highlights the relative success of the attack. Similar conclusions can be drawn regarding hijacking the model for recognizing ethnicity. SnatchML  achieves over $40\%$ accuracy under the white-box setting for the Transformer and ResNet-9 modelsfor a hijacking UB of $75\%$. However, we observe that hijacking gender is more challenging, and the attack has shown its limitations despite its accuracy consistently exceeding random guessing probability.
7. Scenario 3: Pneumonia Diagnosis
7.1. Context

### v2

@@ III SnatchML: General Approach
$p=\underset{i\in[1,m]}{Argmin}\Big{[}\delta(\zeta_{h}(x^{*}_{i}),\zeta_{h}(x_{s}))\Big{]}$ (2)
@@
In essence, SnatchML  takes advantage of the model’s capability to distinguish patterns in a data distribution corresponding to a task $\mathcal{T}^{\prime}$ that was not part of its training, effectively hijacking the model at inference time. SnatchML can be viewed as a form of zero-shot transfer learning, where the victim model, kept entirely frozen, is repurposed for a new unrelated task, in a black-box or white-box setting.
@@
Remark. It is worth noting that in the proposed approach, there are no a priori constraints on $m$. In fact, in the state-of-the-art attacks, such as [38], the number of classes that can be inferred through the hijacking task cannot exceed those in the original task, i.e., $m\leq n$. Therefore, our approach not only assumes a stronger threat model but also overcomes this limitation, enabling the attacker to hijack an $n$-class model for an $m$-class task where $m>n$. We illustrate this case in Section IV. Moreover, since there is no interference with the model during training, the same model can be hijacked to execute more than one task at the same time. This case is illustrated in Section V.
IV Scenario 1: Emotion Recognition
@@
We investigate the vulnerability of a model trained for emotion recognition (ER) to being hijacked at inference time for biometric identification. In this scenario, the adversary has access to a dataset of ’persons of interest’ (the hijacking dataset) and aims to identify these individuals by exploiting the original model.
IV-A Setup
@@
CK+ Dataset. The Extended Cohn-Kanade (CK+) [28] is a dataset for ER that contains $593$ video clips from $123$ individuals between the ages of $18$ and $50$. These videos capture transitions between neutral and peak emotions: anger, disgust, fear, happiness, sadness, and surprise.
@@
Olivetti Faces Dataset. [39] is a facial recognition dataset comprising $400$ face images gathered from 40 unique individuals. The dataset captures variations in lighting and facial expression with a uniform black background, and labeled with the identities of the $40$ individuals.
@@
Celebrity Dataset. We extracted $79$ facial images for $9$ different celebrities. Then, using BRIA AI [8], which is based on a visual generative model [12], we generated emotion-specific images from the neutral images. These images are labeled for seven emotions (anger, disgust, fear, happiness, sadness, surprise, and neutral) and associated with $9$ unique individuals used as identity labels.
@@
Synthetic Dataset. To cover the synthetic data distribution case, we create a new dataset labeled for both emotions and identities using MakeHuman [5], an open-source tool designed to generate virtual human faces. This dataset encompasses $395$ images labeled with both emotion and identity. Among these images, $47$ depict neutral expressions, while the remaining $348$ represent six distinct emotions. The dataset contains a total of $47$ individual identities, each associated with varying emotional expressions.
@@
Figure 2: Examples of the hijacked ER model’s output top-5 candidates for users re-identification from CK+ dataset.
@@
Figure 3: SnatchML performance on hijacking an ER model for users re-identification. ’Hijacking LB’ and ’Hijacking UB’ refer to the lower and upper bounds, respectively.
We evaluate this attack scenario by designing an experimental setup with $4$ pre-trained ER models: (i) 2D-CNN [1, 35] with three convolutional layers, each followed by ReLU activation and max-pooling followed by two fully connected layers, (ii) ResNet-9 [21], (iii) MobileNet [22], and (iv) Vision Transformer [11]. We train the models for ER on CK+ [28] as the original task $\mathcal{T}$ using a learning rate of $0.001$ and Adam optimizer for $100$ epochs.
Comparison. To evaluate the accuracy of the hijacking tasks, we consider a lower bound and an upper bound (UB): The lower bound (LB) corresponds to the random guessing probability, while the UB corresponds to an unconstrained version of [38] without covertness requirement, which is practically equivalent to freely poisoning the model by training it on both the original and hijacking tasks explicitly. Additionally, we compare with the accuracy achieved when the baseline models are trained specifically for the hijacking task, in scenarios where dataset size and the number of samples per class are not limiting factors.
IV-B ER Systems for user re-identification
@@
We refer to this scenario as user re-identification, as it involves the adversary aiming to identify users whose data (images) were used to train the original ER model. It is important to note that in our experiments, the hijacking is performed on users’ data that is not part of the original training dataset, to follow our threat model. We use the top-N ranked reference users as an evaluation metric for the the hijacking attack performance. The top-N metric is particularly relevant for the evaluation of identification tasks [45]; it can provide a contextual method for narrowing down potential candidate users. Figure 2 provides an illustration with queries and their corresponding top-5 samples. In the second line of this figure, the query image features a person with dark skin and only one individual with dark skin appears in the top-5 reference users. The attacker could reasonably deduce that the unknown query image likely belongs to this individual. Figure 3 shows the top-1 to top-5 accuracy results for the hijacking task, presented as mean and standard deviation after conducting $10$-fold experiments. Considering our hijacking task involves identifying $85$ persons, we use random classification as a reference ($1.1\%$). SnatchML achieves over $40\%$ top-1 accuracy for 2D-CNN, while the hijacking UB is up to $98\%$. Although all four tested models demonstrate significant performance, the results shown in Figure 3 reveal disparities among them, suggesting that different architectures may have varying levels of susceptibility to the proposed model hijacking for the same task. For example, while the attack on 2D-CNN and ResNet-9 succeeded with around $40\%$ accuracy, the ViT and MobileNet were more robust with an average Top-5 accuracy not exceeding $30\%$.
IV-C ER systems learn biometric identification
@@
In this second scenario, we consider a setting where the attacker aims to hijack the model for biometric identification of users whose data is not necessarily a part of the ER model’s training dataset. Following the same approach in Section IV-B, we exploit the pre-trained ER model to identify unseen users based on the extracted BEK. As for the hijacking reference database, we consider two main settings:
@@
Case (i): To evaluate the identification attack, we use Olivetti dataset [39]. We perform ER queries with images from Olivetti dataset to get the output FV/Logits and run SnatchML. Figure 4 displays illustration samples of the identification task for Olivetti. An interesting observation is highlighted in the first row, where the top-5 output candidates for a query involving a person wearing glasses also predominantly feature candidates wearing glasses. This indicates that the ER model has learned to recognize this accessory, despite its irrelevance from emotion recognition perspective. The second row in Figure 4 also highlights the significance of the top-5 metric. In this case, the query image corresponds to a female individual with $4$ out of $5$ of the closest candidates representing male subjects, with the first female candidate appearing only at rank-5, potentially leading to a precise identification. Figure 5 shows the hijacking attack accuracy results on the four pre-trained ER models in terms of top-1 to top-5 accuracy. The attacks demonstrate notable success, with the top-1 accuracy achieving over $60\%$ on 2D-CNN and ResNet-9. Although the lower success rate on MobileNet and the ViT remains consistent in this setting, the hijacking performance is still considerable considering a random guessing probability of $\sim 2.5\%$.
Figure 4: ER model used to to identify users from the Olivetti dataset [39]. The top-5 similar users from the hijacking reference database are displayed for each query.
Figure 5: SnatchML performance on hijacking an ER model for biometric identification of users from the Olivetti dataset.
Figure 6: SnatchML Performance in Hijacking an ER Model for Identification Using Real (Celebrity) and Synthetic Datasets.
@@
Case (ii): In this scenario, we assume that the attacker has access only to neutral images of targeted users (e.g. acquired from ID cards). To conduct our attack, we use two datasets featuring individuals with neutral facial expressions, namely Celebrity and Synthetic. The ER model is pretrained on CK+ dataset. We assess the effectiveness of the identification hijacking attack by performing ER queries with images from both Celebrity and Synthetic datasets. The accuracy of the hijacking task using the four pretrained ER models is shown in Figure 6. Notably, the top-1 identification accuracy reached up to approximately $100\%$ in the 2D-CNN model under both black-box and white-box attacks for both datasets. The least successful attack was observed with MobileNet under the black-box scenario, where it achieved approximately $48\%$ top-1 identification accuracy with the Synthetic dataset while the random guess probability is $\sim 2\%$.
V Scenario 2: Age/Gender/Ethinicity
@@
TABLE I: SnatchML Performance on the UTK dataset. We train 04 benign ML models (2D-CNN, ResNet-9, MobileNet, and ViT) to predict Age, Gender, and Ethnicity. We hijack each model to infer other properties (e.g., from age, we infer gender/ethnicity).
@@
Hijacking Age Accuracy Hijacking LB 0.166 0.166 0.166 0.166
@@
Hijacking Gender Accuracy Hijacking LB 0.500 0.500 0.500 0.500
@@
Hijacking Ethnicity Accuracy Hijacking LB 0.200 0.200 0.200 0.200
@@
In this section, we consider an application that predicts personal attributes (e.g., age, gender, ethnicity) from facial images. An example of this application is Microsoft Azure’s Face [31] as an MLaaS API, where users can query one specific model (e.g., for age estimation) using facial images. Given an original task $\mathcal{T}$ (age estimation), an adversary aims to hijack the model to infer other personal attributes like gender or ethnicity. The implications of this hijacking attack can be critical, mainly if users don’t consent to using their personal data. Unauthorized inference of gender and ethnicity from this data could constitute a significant privacy violation, potentially leading to discriminatory practices, especially against individuals from marginalized ethnic or gender groups.
V-A Setup
@@
UTKface Dataset [49]: a collection of $20,000$ facial images of individuals ranging in age from $1$ to $116$ years old. Each image is labeled with information on age (6 classes), gender (2 classes), and ethnicity (5 classes). We consider the same model architectures as in Section IV-A and compare our results with the hijacking upper bound (UB), which corresponds to an unconstrained version of [38] without covertness requirement.
V-B Cross-attribute hijacking: age, gender and ethnicity
@@
The adversary’s goal is to hijack a model – exclusively trained on one of the three tasks– to infer other personal attributes for which the victim model has not been trained. Specifically, from an unknown facial image query submitted by an unknown user to an age estimation model and by only using the original model’s BEK, the adversary aims to infer the gender and ethnicity of the user. For a comprehensive study, we test all the possible combinations of (age, gender, and ethnicity) as original and hijacking tasks.
@@
Table I provides an overview on the experimental results with random guess probability as a lower bound. Different tasks have varying levels of difficulty that can be related to their inherent complexity. For example, age recognition is a challenging task; a 17-year-old ’teenager’ and an 18-year-old ’young adult’ are not easily distinguished. The fact that models trained for age prediction as the original task achieve accuracy ranging from $63\%$ (Transformer) to $70\%$ (ResNet-9) illustrates the task complexity and highlights the relative success of the attack. Similar conclusions can be drawn regarding hijacking the model for recognizing ethnicity. SnatchML achieves over $40\%$ accuracy under the white-box setting for the Transformer and ResNet-9 models for a hijacking UB of $75\%$. However, we observe that hijacking ’Gender’ is more challenging, and the attack has shown its limitations despite its accuracy consistently exceeding random guessing probability.
VI Scenario 3: Pneumonia Diagnosis